OpenStudy (anonymous):

Graph the system of constraints. Name all vertices. Then find the value of x and y that maximize or minimize the objective function. Find the maximum or minimum value. 3x+y<_7 x+2<_9 x>_0, y>_0 Maximum for P= 2x+y

OpenStudy (ybarrap):

$$ 3x+y <7\implies y < 7-3x\\ x+2<9\implies x < 7\\ x>0,y>0\implies\text{ Range and Domain is in Quadrant I}\\ \therefore 0<y<7-3x\text{ & } 0<x<7 $$ http://www.wolframalpha.com/input/?i=0%3Cy%3C7-3x%2C+0%3Cx%3C7 The vertices are easily found: 1) \(x=7\) and \(y=0\) 2)\(x=0\) and \(y=\cfrac{7}{3}\) 3) \(x=0\) and \(y=0\) Using these three vertices, plug in to : \(P=2x+y\) and determine which of these <x,y> pairs makes P the largest.

OpenStudy (ybarrap):

*The vertices are easily found: $$ 1) ~x=\cfrac{7}{3},y=0\\ 2) ~x=0,y=0\\ 3) ~x=0,y=7 $$

OpenStudy (anonymous):

is this the answer?

OpenStudy (ybarrap):

The answer is 1, 2 or 3 above. Plug in <x,y> into 2x+y for each and select the <x,y> that makes P the largest. For example, for 1): x=7/3,y=0, so P=2x+y=2(7/3)+0=14/3. This is P at <7/3,0>. Next, find P at <0,0> and at <0,7> and compare to P at <7/3,0>, which we just found to be 14/3.

OpenStudy (anonymous):

BTW the <_ is actually